OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = −three fourthsx + 8 y = −three fourthsx + 3

OpenStudy (souvik):

what is the slope of the equation?

OpenStudy (anonymous):

is it (-3,4)

OpenStudy (souvik):

okay y=?

OpenStudy (anonymous):

y=4/3

OpenStudy (souvik):

no...

OpenStudy (anonymous):

idk then what is it??

OpenStudy (souvik):

4x-3y=-12 or 3y=4x+12 of y=4/3x+4 so the slope is 4/3 get it?

OpenStudy (anonymous):

wait is that the answer?

OpenStudy (souvik):

no ...

OpenStudy (anonymous):

so what is the answer?

OpenStudy (souvik):

y-4=4/3(x+3) or y=4/3x+8

OpenStudy (anonymous):

thanks!!

OpenStudy (souvik):

do u understand?

OpenStudy (anonymous):

kinda i have another question?

OpenStudy (souvik):

ok

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (souvik):

find the slope first...

OpenStudy (anonymous):

the slpoe is (6,-1) right??

OpenStudy (souvik):

no...

OpenStudy (anonymous):

is it -3,7

OpenStudy (souvik):

3x+y=7 or y=-3x+7 so the slope is -3

OpenStudy (anonymous):

oh so okay what do i do after that ??

OpenStudy (theeric):

Hi! Let me catch up, and then I'll be able to help!

OpenStudy (theeric):

Hi! First I'll lay out how to solve this problem. You've already made progress. You want to find a line perpendicular to the original line. By definition, you want a line that has a slope that's a reciprocal of the original line's.

OpenStudy (theeric):

So you found the original line's slope with souvik's help. Do you know what the slope is?